# Elimination Without Eliminating: Computing Complements of Real Hypersurfaces Using Pseudo-Witness Sets

### v1

@@ Acknowledgments
PB was supported by DFG, German Research Foundation – Projektnummer 445466444. JC acknowledges the support of the National Science Foundation Grant DMS-2402199. AKE was partially supported by National Science Foundation Grant DMS-2023239. JDH was partially supported by the National Science Foundation Grant CCF-2331400, Simons Foundation SFM-00005696, and the Robert and Sara Lumpkins Collegiate Professorship. OH was partially funded by the European Union (Grant Agreement no. 101044561, POSALG). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Research Council (ERC). Neither the European Union nor ERC can be held responsible for them. JLG was partially supported by the Robert and Sara Lumpkins Collegiate Professorship.

### v2

@@ Acknowledgments
PB was supported by DFG, German Research Foundation – Projektnummer 445466444. JC acknowledges the support of the National Science Foundation Grant DMS-2402199. AKE was partially supported by National Science Foundation Grant DMS-2023239. JDH was partially supported by the National Science Foundation Grant CCF-2331400, Simons Foundation SFM-00005696, and the Robert and Sara Lumpkins Collegiate Professorship. OH was partially funded by the European Union (Grant Agreement no. 101044561, POSALG). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Research Council (ERC). Neither the European Union nor ERC can be held responsible for them. DKJ and DM were partially supported by NSERC Discovery Grant RGPIN-2023-03551. JLG was partially supported by the Robert and Sara Lumpkins Collegiate Professorship.